OpenStudy (anonymous):

Question 7.7. The weights of bags of ready-to-eat salad are normally distributed with a mean of 290 grams and a standard deviation of 10 grams. What percent of the bags weigh less than 280 grams? (Points : 3) 34% 13.5% 16% 36.5%

OpenStudy (anonymous):

\[\begin{align*} P(X<280)&=P\left(\frac{X-290}{10}<\frac{280-290}{10}\right)\\ &=P(Z<-1) \end{align*}\] You can determine this probability using a table or http://en.wikipedia.org/wiki/68%E2%80%9395%E2%80%9399.7_rule

OpenStudy (anonymous):

@hartnn